OpenStudy (anonymous):

After 5 coinflips, what is the probability that there are 3 heads?

OpenStudy (anonymous):

I think it is \[\frac{5!}{3!(5-3)!}\times(\frac{1}{2})^5\]but why?

OpenStudy (anonymous):

so there are \[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\]combinations where I get 3 heads, and for each of those combinations the probability is 0.5^5 I get it now :-D

OpenStudy (jamesj):

Write down outcomes of five coin tosses like this: five tails: ttttt 2 heads, 2 tails then a head: hhtth etc. Then the probability of 3 heads is the probability of any of the possible outcomes with 3 heads: that is, hhhtt, hhtht, hhtth, htthh, ... Now how many outcomes are there like this? There are 5 choose 3, 5C3 = 5!/(3! (5-3)!) = 5!/(3! 2!)

OpenStudy (jamesj):

and each of those has probability (1/2)^5

OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

you're like the next Srinivasa Ramanujan

OpenStudy (jamesj):

The othe rway to think about this let H = probability of a head, T = probability of a tail. Then (H + T)^5 gives all the possibilities of # heads and tails and their probabilities, and we want the term H^3.T^2. By the binominal theorem, it has coefficient 5C3 hence the probability of 3 heads is \[{5 \choose 3} H^3 T^2 = {5 \choose 3} (1/2)^5\]

OpenStudy (jamesj):

This might seem slightly over the top notation for just heads and tails, but it's very helpful for other problems. For example: suppose we throw a fair six-sided die seven times. What's the probability we get 5 or 6 three times. Well the outcome we want, 5 or 6, has probability P = 1/3, and the other outcome has probability Q = 2/3 Hence the probability of three 5 or 6 is\[{7 \choose 3} P^3Q^4 = {7 \choose 3} (1/3)^3 (2/3)^4\]

OpenStudy (anonymous):

Great stuff. I feel enlightened :-D